Hey! I have a question on how to solve this. I have to use geometric/arithmetic sequences in order to solve this word problem.
A farmer gathers 35 bushels of potatoes on the first day of harvest. The farmer estimates that on each successive day of the harvest, the amount gathered can be 4 bushels more that the preceding day. If the harvest lasts 14 days, what is the total number of bushels harvested?
Thanks! :D

Answer by mananth(16949)

A farmer gathers 35 bushels of potatoes on the first day of harvest. The farmer estimates that on each successive day of the harvest, the amount gathered can be 4 bushels more that the preceding day. If the harvest lasts 14 days, what is the total number of bushels harvested?

we have to calculate total harvest for 14 days
first day = 35 = a
next day = 35+4 so d= 4
number of days = 14 =n
sn=n/2{2a+(n-1)d]
plug
Sn = 14/2[2*35+(14-1)*4]
=7(70+52)
=7*122
=854 bushels